lexxxxxxxxx:

f(x)=log(x+3) inverse?

Vocaloid:

replace f(x) with y y = log(x+3) switch x and y to get x = log(y+3) ^ try solving this equation for y, that will give you the inverse function

lexxxxxxxxx:

how will i solve the equation for y ?

Vocaloid:

so you have log(y+3) on the right side, to "undo" a logarithm, raise both sides to the base of 10 so you get 10^x = y + 3, then you just need to subtract 3 from both sides to get y

lexxxxxxxxx:

7^x=y is what i got

Vocaloid:

10^x and 3 are not like terms so you cannot subtract them to get 7^x 10^x = y + 3 subtracting 3 from both sides gives us y = 10^x - 3 as your inverse

Vocaloid:

using the original notation \(f^{-1}(x)\)= 10^x - 3

lexxxxxxxxx:

yes i was about to say that ! thanks again.